Write the equation of each conic section with the given properties:

A hyperbola with foci at (15,0) and (-15,0) and asymptotes y = 3/4x and y = -3/4x


1
Expert's answer
2021-02-24T06:49:04-0500

Solution:

"\dfrac{x^2}{a^2}-\dfrac{y^2}{b^2}=1"


The info about the asymptotes indicates that "\frac{b}{a}" = "\frac{3}{4}".

The distance from either focus to either asymptote is "b,"

so I examine the line through (15,0), perpendicular to y = ("\frac{3}{4}")x.

Its equation will be y = -("\frac{3}{4}")(x - 15) = 20 - ("\frac{4}{3}")x. It intersects y = ("\frac{3}{4}" )x where 20 = ("\frac{25}{12}")x, or x = "\frac{48}{5}" , y = "\frac{36}{5}" .

The distance from that point to (15,0) is

sqrt[("\frac{27}{5}")2 + ("\frac{36}{5}")2] = 3*"\sqrt{(\frac{3}{5})^2+(\frac{4}{5})^2}=3".


So the equation of the hyperbola is


"\dfrac{x^2}{4}-\dfrac{y^2}{3}=1" .